answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of drawing one joker from a deck of 52 cards?

The question should be ..............from a deck of '54' cards.

P(a joker from 54 cards) = 2/54 = 1/27

How do you determine legitimate probability distributions?

Well... the probabilities should add up to exactly 1 and cannot be negative.

What is the the relationship between population and sample parameter and statistic?

The relations depend on what measures.

The sample mean is an unbiased estimate for the population mean, with maximum likelihood.

The sample maximum is a lower bound for the population maximum.

How do you refill ace liner 502?

This isn't filled like your normal standard stapler. First, pull up the top half of the stapler (the actual part that holds the staples and basically does all the work. Second, push in the two side buttons that are about 1/4 of the way down from the front of the stapling head. This will open the staplereservoir. There is a rod inside that has a round disc that comes out the back. Grab ahold of the disc and push in then lift up. This will release the rod. Pull the rod all the way back, as far as it will go. It doesn't come out, it will stop once it reaches the farthest back position. Lift the rod up to approximately a 90 degree angle. The staples go inside the opening that is now exposed. The basically rest along a rail. Push them in as far as they go and then return the rod you just pulled out and lifted up to it's natural parallel position. Push the rod back in, allowing the spring on the rod to create the tension needed to hold the staples in place and snap the catch clip near the back of the rod (just in front of the disc) back down into place. That's it! It sounds more complicated than it is but it's really not at all once you do it one time.

Probability of picking one red card of a deck of 52 playing cards?

The probability of picking one red card of a deck of 52 playing cards is 26 out of 52, or 1 out of 2.

What is probability of picking a King from a deck of cards?

There are 4 Kings in a deck of 52 cards, so the probability of picking a King is 4/52 or 1/13.

The Associative Property does not work for which operations?

Subtraction and division.

While 2+ (3+4) = 2+ (4+3), the subtraction 2-(3-4) ≠ 2-(4-3). One yields 3 while the other yields 1.

Similarly, multiplication has this property while division does not.

How many possible outcomes if you roll a die 4 times?

It depends what you mean by outcome. If you are interested in the total number of different outcomes, then there are 6 x 6 x 6 x 6 or 1296 unique possible outcomes. If you are only interested in the total (the sum) of the resulting roll, it can be anything from 4 to 24 so there are just 21 possible outcomes.

Is the Z-test the same as Z-scores?

No, the Z-test is not the same as a Z-score. The Z-test is where you take the Z-score and compare it to a critical value to determine if the null hypothesis will be rejected or fail to be rejected.

An event that will definitely happen is an event with what percent probability?

An event that will definitely happen is an event with 100% probability.

If a coin is flipped twice what is the probability that both will be heads?

The sample space is HH, HT, TH, HH. Since the HH combination can occur once out of four times, the probability that if a coin is flipped twice the probability that both will be heads is 1/4 or 0.25.

Why does a researcher want to go from a normal distribution to a standard normal distribution?

A researcher wants to go from a normal distribution to a standard normal distribution because the latter allows him/her to make the correspondence between the area and the probability. Though events in the real world rarely follow a standard normal distribution, z-scores are convenient calculations of area that can be used with any/all normal distributions. Meaning: once a researcher has translated raw data into a standard normal distribution (z-score), he/she can then find its associated probability.

What is the probability of drawing a spade from a deck of cards?

A standard deck is 52 cards. These 52 cards are divided evenly into 4 suits - spades, clubs, diamonds, and hearts. There are 13 cards in each suit, ace through king. Therefore, there are 13 spades in a deck of 52 cards. 13 is one fourth of 52 which makes the probability of drawing a spade 1/4, .25, or 25 percent.

Two coins are tossed. What is the probability of tossing at least one head?

There are 2 coins, 1 and 2. Each has two possibilities, H or T.

The possibilities are:

1H, 2H

1H, 2T

1T, 2H

1T, 2T

Each possibility has an equal chance of happening.

The chance of tossing at least one head is 3/4.

Two unbiased dice are thrown Find the probability that neither double nor a total of 10 will appear?

The sample space is 62 or 36. The easier way to do this problem is to find the probability of doubles or 10 and subtract from 1. The probability of a double is 6/36 (1,1; 2,2; 3,3; 4,4; 5,5; 6,6). There are 3 ways to get a 10; 4,6; 6,4; & 5,5. We already counted 5,5 in the double so there are 2/36 we add to the 6/36 to obtain 8/36. So, 1-8/36 = 28/36 or 7/9 or 0.7778 probability that neither double nor a total of 10 will appear.

A bag contains 6 black and 3 white balls Another bag contains 5 black and 4 white balls If one ball is drawn from each bag find the probability that these two balls are of the same color?

14/27 ~ 0.519 ~ 51.9%


----------------------------------------------------------------------------------------------------

EXPLANATION

For bag 1: Probability of drawing a black ball; P(B1) =

6/9. A white one; P(W1) =3/9

For bag 2: P(B2) =

5/9. P(W2) =

4/9.

Probability of drawing a black ball from each bag: P(B1UB2) =

(6/9)∙(5/9) =

10/27

Probability of drawing a white ball from each bag: P(W1UW2) =

(3/9)∙(4/9) =

4/27

The probability of drawing 2 balls of the same color is the sum of the above

probabilities:

P(2 same color balls) =

P(B1UB2) + P(W1UW2) =

10/27 +4/27 =

14/27 =

0.5185...

P(2 same color balls) ~ 0.519 ~ 51.9%

Trending Questions
If you flip 10 pennies all at the same time what is the probability that exactly 3 will come up heads and 7 will come up tail? What is the probability of flipping a coin twice and it landing on heads exactly one time? What is the total number of possible outcomes that the number of different ways that first and second prizes can be determined from a total of 10 participants? What does is mean when the standard deviation is 0? Exactly three people from a line at the grocery store what is the probability that they will line up in descending order of age oldest middle and youngest? What is absolute sample size? How does a motor function? What is the equation used to determine probability in what ways can probability be expressed? How many 3's will i get if i roll a dice 100 times? What is the probability of drawing a blue card replacing it and then drawing a blue card? How many number 7 cards are in a deck of 52 cards? Compute the probability of rolling seven with two six sided dice? When a random sample of 64 bottles of the 200 count Bayer Aspirin showed a mean turnover rate of 5.84 Shall you reject the hypothesis that the population mean is 6.0? How are populations and samples symbolized? A complete probability distribution is always an objective listing of all possible events Since it is impossible to list all the possible outcomes from a single event probability distributions are o? If you draw a card from a standard deck of cards what is the probability of drawing a nine? If a coin is tossed and then a dice is rolled what is the probability of getting a 5 on the dice given that the coin landed tails up? Ron decided to use blocking to deal with extraneous factors in his experiment. If 4 treatments are applied to one of his groups how many treatments should be applied to each of his other groups? What is the probability of rolling a pair of dice and obtaining a total score of 9 or more? Where is the percent sign on the keyboard?